OpenStudy (anonymous):

Is the square root of (9/4) end up being: (3/2)?

OpenStudy (anonymous):

*Does

OpenStudy (anonymous):

You can rewrite \[\sqrt{\frac{9}{4}} = \frac{\sqrt{9}}{\sqrt{4}} = \frac{3}{2}\]:-)

OpenStudy (anonymous):

So yeah, it does :-)

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

No problem.